a person travels two parts of the total distance in the ratio 2 is to 1 with constant constant speed of 30 kilometre per hour and 40 km per hour respectively what is the average speed of the journey
1
Expert's answer
2018-04-24T08:55:08-0400
The average speed of an object is defined as the total distance travelled by the object by total time taken by the object to cover the total distance.

average speed = total distance / total time taken

Given the first part of the total distance is twice that of the second part.
Let us assume that first part is 240 km and second part is 120 km. The ratio of two parts is 2:1 .

The speed of the person when travelling through the first part = v1 = 30 km/h
The time taken to cover the first part = t1 = 8 hours

The speed of the person when travelling through the second part = v2 = 40 km/h
The time taken to cover the second part = t2 = 3 hours

Therefore,
Total distance = 240 + 120 = 360 km
Total time taken = 8 + 3 = 11 hours

Thus,
Average speed = 360 / 11 = 32.72 km/h